Javaid is much younger to Amir as he is older to Tariq, if the sum of the ages of Amir and Tariq is 48 years, what is the age of Jawaid??
Answer by ankor@dixie-net.com(22740)

Javaid is much younger to Amir as he is older to Tariq, if the sum of the ages of Amir and Tariq is 48 years, what is the age of Jawaid??
:
The ages in ascending order: T---J---A
:
J - T = A - J
J + J = A + T
We know A + T = 48, therefore
2J = 48
J = 24 yrs old